Which is the best approximation of the solution of the equation:
-2.9n - 4.89 = 1.75(n + 14.5)?
A. -17.81 B. -6.51
C. -4.55 D. -4.41

Answer by checkley75(3666)

2.9 becomes 3
4,89 becomes 5
1.75 becomes 2
14.5 becomes 15
thus:
-3-5=2(n+15)
-8=2n+30
2n=-8-30
2n=-38
n=-38/2
n=-19 answer
closest answer of choice is A)